Data structure and algorithms

6112 pointsj One of the importast operations on a RST is to find sway to in the tree. Considering the folleowing binary search tre trave 15 45 (a) Please give the in-order traversal: hi Please gee the pre-onder traversal e) Plrase ge the post-order traversal
media%2F65e%2F65e1b0fc-4d3d-4d33-a34e-28
Answer #1

a- logn + 10000 - answer is O(logn n)
b-nlog n + 15n + 0.0002n^2 answer is O(n^2)
c-100000n^2 + 16n + 2^n answer is O(n^2)
d-n + (n-1) + (n-2)+ - - - - 3 + 2 +1 = n(n+1)/2 which is n^2 +n/2 so answer is O(n^2)
e- 2^10 + 3^5 = O(1)

13 16 68 34 the image in the first image is not clear so cannot provide the tree traversals, if you can provide a clear image I can provide the answer for those as well.